# ASVD: Activation-aware Singular Value Decomposition for Compressing Large Language Models

### v1

@@ Abstract
This paper explores a new post-hoc training-free compression paradigm for compressing Large Language Models (LLMs) to facilitate their wider adoption in various computing environments. We delve into the challenges of LLM compression, notably their dependency on extensive training data and computational resources. We propose a training-free approach dubbed Activation-aware Singular Value Decomposition (ASVD) to address these limitations. ASVD effectively manages activation outliers by adjusting the weight matrix based on the activation distribution, improving decomposition accuracy and efficiency. Our method also addresses the varying sensitivity of different LLM layers to decomposition, with an iterative calibration process for optimal layer-specific decomposition. Experiments demonstrate that ASVD can compress network by 10%-20% without losing reasoning capacities. Additionally, it can be seamlessly integrated with other LLM compression paradigms, showcasing its flexible compatibility. Code and compressed models are available at ASVD4LLM.
@@ 1 Introduction
Figure 1: Our post-training LLM decomposition method is orthogonal to existing LLM compression techniques, enabling it to function as a versatile and plug-and-play solution for prevalent compression paradigms, including popular quantization methods.
@@
Numerous strategies have been proposed to mitigate the memory consumption issues in Large Language Models (LLMs), as surveyed by Zhu et al. [2023]. These methods predominantly fall into two categories: neural network compression and system optimizations. In terms of neural network compression, techniques such as weight quantization [Dettmers et al., 2022, Shang et al., 2023], network pruning [Frantar & Alistarh, 2023], and knowledge distillation [Agarwal et al., 2023] have been extensively explored. Meanwhile, system optimizations focus on efficient memory management, with cache mechanisms like FlashAttention [Dao et al., 2022] and PagedAttention [Kwon et al., 2023]. Distinct from these approaches, the paradigm of low-rank decomposition is less explored but promising. This technique involves approximating the weight matrices in neural networks with matrices of lower rank, effectively reducing the model size. Given the massive number of parameters in LLMs, low-rank decomposition offers significant potential for memory reduction. Despite its potential, it remains a relatively underutilized approach in the field of LLM compression. Furthermore, low-rank decomposition is complementary to existing LLM compression techniques by further compressing quantized or pruned models, enhancing overall efficiency [Cheng et al., 2017, Deng et al., 2020].
@@
From the perspective of network compression, traditional low-rank decomposition methods typically adhere to a straightforward process: initially training the original model and subsequently fine-tuning the decomposed model [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021, Hsu et al., 2022]. While this approach is effective, it is resource-intensive and requires the entire training dataset and substantial computational power for end-to-end backpropagation. Applying this method to Large Language Models (LLMs) would encounter major challenges. Firstly, the training data for LLMs may not always be readily available, often restricted by privacy and commercial considerations. Secondly, the training process for these models is notoriously expensive, both in terms of time and computational resources. Given these constraints, the concept of “training-free” compression emerges as a more viable approach for LLMs [Zhu et al., 2023]. This approach includes methods like LLM post-training quantization[Dettmers et al., 2022, Yuan et al., 2023] and LLM post-training pruning [Frantar & Alistarh, 2023], which compress LLMs without the need for extensive retraining. These training-free methods offer a more practical solution for efficiently compressing LLMs.
@@
To realize low-rank post-training decomposition in LLMs, we conduct an extensive analysis of the baseline methods for LLM decomposition. We first observe that straightforward application of existing low-rank decomposition techniques, which typically necessitate training, proves ineffective for LLMs [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015, Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021]. Digging into the failures, we reveal several challenges unique to post-training decomposition for LLMs. One key challenge involves managing outliers in the activations, which significantly intensifies the decomposition error. This challenge mirrors findings in LLM quantization research [Lin et al., 2023, Kim et al., 2023], highlighting the importance of handling such outliers in LLMs. To address these challenges, we introduce Activation-aware Singular Value Decomposition (ASVD). Specifically, ASVD incorporates the distribution of activations into the decomposition process. It achieves this by scaling the values in the weight matrix column-wisely, based on the distribution patterns observed across input and output channels. This adjustment proves particularly beneficial for channels with outlier activations, allowing the ASVD to allocate enhanced focus to these specific weights. Such targeted adjustment facilitates a more accurate reconstruction of the weight matrix, thereby minimizing prediction errors and improving the overall efficiency of the decomposition process.
Our research further delves into the varying sensitivity of different LLM layers to decomposition. We discovered that weights in Multi-Head Attention (MHA) layers [Vaswani et al., 2017] tend to be more resilient to decomposition compared to those in Multi-Layer Perceptron (MLP) layers. This sensitivity is not uniform across all layers, prompting us to develop an method to assign the compression ratio for each layer. This method assesses each layer’s sensitivity to decomposition at different ranks, enabling us to assign a suitable rank for optimal decomposition. This process is notably efficient, requiring only a limited sample set for evaluation. Additionally, we optimize inference by absorbing singular values after decomposition, resulting in computationally efficient matrix operations. Our approach evenly distributes singular values into other U and V matrices, benefiting weight quantization by reducing quantization errors.
We applied our method to open-access LLMs, specifically LLaMA [Touvron et al., 2023a] and LLaMA-2 [Touvron et al., 2023b]. Our experiments reveal that ASVDcan effectively compress network weights by approximately 10%-20%, with only a marginal 1% accuracy loss on benchmarks like MMLU [Hendrycks et al., 2020]. This compression is achieved in a training-free manner, exemplified by the decomposition process for LLaMA-7b, which takes merely 4 hours on an Nvidia A6000 GPU. Importantly, we conducted preliminary validations of ASVD compatibility with 4/6/8-bit quantization. ASVD emerges as an orthogonal paradigm for LLM compression, seamlessly integrating as a plug-and-play technique that aligns with existing LLM quantization methods.
@@ 2.1 Large Language Model Compression
The field of model compression for Large Language Models (LLMs) has seen a surge of innovative techniques aimed at mitigating the substantial computation and memory requirements these models demand [Zhu et al., 2023]. Various methods have emerged to address this challenge, each taking a unique approach to reduce the memory footprint of LLMs. These methods primarily fall into three categories: weight quantization [Courbariaux et al., 2015, Dettmers et al., 2022], network pruning [LeCun et al., 1989, Frantar & Alistarh, 2023], and knowledge distillation [Hinton et al., 2015, Agarwal et al., 2023]. For the wide body of research on LLM compression, please refer to [Zhu et al., 2023] for the comprehensive survey. Among these methods, weight quantization has gained significant traction in the context of LLMs due to its effectiveness. However, despite its popularity as a neural network compression technique, low-rank factorization has not been extensively explored in the realm of LLMs. Recognizing this gap, we introduce a novel low-rank decomposition method tailored specifically for decomposing the weight matrices of LLMs in a training-free manner. This approach aims to provide an efficient and effective solution for LLM compression, leveraging the untapped potential of low-rank factorization in this domain.
2.2 Low-rank Decomposition
The use of low-rank matrix decomposition as a technique for compressing Deep Neural Networks (DNNs) represents a straightforward yet effective approach, garnering considerable attention within both scientific computing and machine learning domains. In recent years, the challenge of efficiently compressing and accelerating large-scale neural networks via low-rank methods has become a focal point of research. This has led to significant advancements in developing and refining low-rank factorization strategies tailored for DNNs [Schotthöfer et al., 2022].
@@ 2.2 Low-rank Decomposition
In the realm of low-rank decomposition for neural network compression, existing methods can be broadly classified into two categories: fixed low rank and variable low rank approaches. Fixed rank methods typically involve decomposing weight matrices of pre-trained networks using techniques like Singular Value Decomposition (SVD) or tensor decomposition, followed by fine-tuning the factorized network [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015]. They also involve constraining weight matrices to maintain a fixed low rank during training [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021], or constructing layers as linear combinations of layers with varying ranks [Ioannou et al., 2015]. A notable limitation of these methods is the introduction of matrix decomposition rank as an additional hyperparameter requiring fine-tuning. In contrast, rank-adaptive methods address this limitation by automatically determining and adjusting the low-rank structure. In particular, Kim et al. [2015, 2019] apply heuristics search to pre-determine the decomposition rank, while Wen et al. [2017] learn low-rank weights through a loss function penalizing approximated matrix ranks. These approaches remove the need for selecting hyper-parameters of LLMs and highlight the potential of low-rank decomposition as an effective tool for model compression, paving the way towards more efficient and scalable Large Language Models (LLMs).
@@
In this work, we introduce our ASVD, a post-training LLM decomposition approach enabling the adaptive determination of SVD ranks to optimize the matrix approximations based on feature activations.
Difference with TensorGPT [Xu et al., 2023]. In the content of LLM compression via decomposition, the most related work is the concurrent TensorGPT [Xu et al., 2023, Zhu et al., 2023], in which the embedding layer of LLMs is compressed through Tensor-Train Decomposition (TTD) [Oseledets, 2011] in order to store large embeddings in a low-rank tensor format, with much fewer parameters. However, there are several differences between those two methods: (1) Unlike TensorGPT which focuses solely on the token embedding matrix, ASVDaims to compress the entire weight spectrum of LLMs. This holistic approach addresses a more critical aspect of LLM compression, as highlighted in recent studies [Lin et al., 2023, Kim et al., 2023]; (2) From the perspective of low-rank decomposition categorization, our method can realize the low-rank decomposition in a rank-adaptive manner, contrasting with the fixed or predetermined ranks used in TensorGPT.
@@ 3 Method
In this section, we initially outline the general process of implementing Singular Value Decomposition (SVD) for weight matrix compression in LLMs, as detailed in Section 3.1. Following this, Section 3.2 identifies and discusses two major challenges associated with implementing SVD-based methods in LLMs. To address these challenges, we propose our Activation-aware Singular Value Decomposition (ASVD) and Sensitivity-aware Truncation Rank Searching (STRC) in Sections 3.3 and 3.4, respectively.
@@ 3.1 Naïve SVD for Compressing Weight Matrix
A large language model (LLM) is primarily based on the Transformer model [Vaswani et al., 2017]. A significant amount of memory and computational overhead is incurred by Transformers’ linear weight matrices, especially in its multi-head attention (MHA) and feed-forward layers [Tay et al., 2022]. Consequently, we focus primarily on decomposing these linear weight matrices in this paper. Singular Value Decomposition (SVD) is employed for this purpose, which involves decomposing a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ into three matrices: $\mathbf{U}$, $\mathbf{\Sigma}$, and $\mathbf{V}^{T}$, such that $\mathbf{W}\approx\mathbf{W}\mathbf{\Sigma}\mathbf{V}^{T}$), where $\mathbf{\Sigma}$ is an $m\times n$ diagonal matrix, the diagonal values in $\Sigma$ are the singular values of $\mathbf{W}$, and $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are corresponding right and left singular vector matrices, respectively [Demmel, 1997].
The SVD compression process for a weight matrix can be summarized in three steps: Decomposition: Factorize $\mathbf{W}$ using SVD. Truncation: Retain the top $k$ singular values and their corresponding right and left singular vectors. This results in approximated matrices $\mathbf{U}_{k}$, $\mathbf{\Sigma}_{k}$, and $\mathbf{V}_{k}^{T}$, where the right singular vector matrix $\mathbf{U}_{k}$ is $m\times k$, singular $\mathbf{\Sigma}_{k}$ is $k\times k$, and left singular vector matrix $\mathbf{V}_{k}^{T}$ is $k\times n$. The choice of $k$ is critical in balancing the compression ratio and the compressed model’s performance. Reconstruction: Reconstruct an approximated weight matrix:
$\mathbf{W}_{k}=\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{U}_{k}^{T}.$ (1)
The primary factor impacting the performance of a model in the context of SVD is the truncation error. This error arises due to the approximation of truncated SVD, which can be measured by calculating the difference between the original weight matrix $\mathbf{W}$ and the approximated matrix $\mathbf{W}_{k}$ as:
$\mathcal{L}_{\text{t}}=\|\mathbf{W}-\mathbf{W}_{k}\|_{F}.$ (2)
where $\|\|_{F}$ denotes the Frobenius norm. Exploiting the properties of SVD, we can further refine this error in terms of the singular values not included in the truncated matrix $\mathbf{W}_{k}$. Assuming the singular values in $\mathbf{\Sigma}$ are arranged in decreasing order, the truncation error can be described as the square root of the sum of the squares of the omitted singular values. This is also why truncated SVD is understood as the best low-rank approximation as it preserves the information most compactly [Eckart & Young, 1936]. Specifically, it is given by: $\mathcal{L}_{\text{t}}=\sum_{i=k+1}^{\min(m,n)}\delta_{i}^{2}$, where $\delta_{i}$ are the singular values that are not included in $\mathbf{\Sigma}_{k}$.
The choice of rank $k$ in SVD plays a crucial role in balancing the compression ratio and the resulting error. Specifically, opting for a smaller rank $k$ typically leads to a higher compression ratio but at the cost of increased truncation error and vice versa. However, the rate of this increase depends on the distribution of the singular values in the original matrix $\mathbf{W}$. If $\mathbf{W}$ possesses a few large singular values while the rest are considerably smaller, a low-rank approximation can effectively capture most of the matrix’s patterns. In such scenarios, even a relatively small $k$ can result in a modest truncation error. Conversely, in cases where the singular values of $\mathbf{W}$ do not exhibit rapid decay, achieving a low truncation error necessitates opting for a higher rank $k$. This variance in singular value distribution underscores the need for careful selection of $k$ to optimize both compression and model performance in the context of LLM decomposition.
@@ 3.2 Challenges of Compressing LLMs via SVD
The concept of decomposing the large matrices in LLMs (e.g., $4096\times 4096$ matrices are ubiquitous in LLaMA-7b [Touvron et al., 2023a]) into lower ranks presents a viable pathway for model compression. However, straightforward application of existing low-rank decomposition techniques, which typically necessitate training, proves ineffective for LLMs [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015, Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021]. The primary limitation of these methods is their reliance on training data and a large amount of optimization for rehabbing the decomposed models, contrasting with our objective to compress LLMs in a training-free manner. To address this gap, we conducted a comprehensive analysis to understand the failure points of traditional methods (i.e., SVD for large matrix in LLM) when applied to LLM compression. Our investigation revealed two primary challenges specific to post-training decomposition of LLMs.
Figure 2: Comparison between SVD and Activation-aware Singular Value Decomposition (ASVD). Outlier channels in input activations ($\mathbf{X}$) are highlight in red, and ASVD takes these into consideration, which can contribute to a reduction in output error.
Challenge 1: Sensitivity of Weight Variation: Recent studies in LLM post-training quantization [Lin et al., 2023, Wei et al., 2022, 2023, Frantar et al., 2022] have underscored the importance of considering activations when compressing LLM weights. This perspective shifts the focus from solely relying on the truncation error $\mathbf{L}_{\text{t}}$, which depends only on the model’s weights, to also accounting for the activations (see Eq.2). The rationale behind this is the critical role of outliers in activations within LLMs  [Lin et al., 2023, Wei et al., 2022, Kim et al., 2023]. Thus, for effective LLM decomposition, our objective optimization becomes:
@@
To illustrate this concept, we visualize an example of $\mathbf{W}$, $\mathbf{W}_{k}$ (decomposed by simply SVD), $\mathbf{X}$, and the resulting variation in activations $\Delta\mathbf{Y}$ in Fig.2 (Top line). This visualization reveals a critical insight: even when the variation in weights $\Delta\mathbf{W}=\mathbf{W}-\mathbf{W}_{k}$ is relatively minor, the corresponding variation in activations $\Delta\mathbf{Y}$ can be huge. This significant variation in activations is a key factor why a straightforward SVD-based decomposition approach falls short in effectively decomposing LLMs. The activation variations, despite being derived from input activations large magnitude (not the weight variations), can lead to considerable changes in the whole model’s output, thereby undermining the decomposition’s efficacy.
@@
These challenges underscore the necessity for innovative approaches specifically designed for the LLM decomposition. Our objective is to achieve efficient compression while circumventing the substantial computational and data demands associated with training-based methods. To address the first challenge, we introduce an Activation-aware SVD mechanism, which is detailed in Section 3.3. This method is designed to mitigate the impact of weight variation on activations. For the second challenge, we propose a Sensitivity-based Truncation Rank Searching mechanism, elaborated in Section 3.4, which adapts to the varying singular value distributions among different layers. These tailored strategies aim to ensure that the decompression maintains the functional integrity of the LLMs while optimizing the memory and computational efficiency.
3.3 Activation-aware Singular Value Decomposition
Our ASVD is a technique designed to refine the weight matrix $\mathbf{W}$ in LLMs by taking into account the influence of input activation channels. The process comprises the following three steps:

### v3

@@ Abstract
In this paper, we introduce a new post-training compression paradigm for Large Language Models (LLMs) to facilitate their wider adoption. We delve into LLM weight low-rank factorization, and find that the challenges of this task stem from the outlier phenomenon in the LLM activations and the sensitivity difference among various kinds of layers. To address these issues, we propose a training-free approach called Activation-aware Singular Value Decomposition (ASVD). Specifically, ASVD manages activation outliers by scaling the weight matrix based on the activation distribution, thereby enhancing decomposition accuracy. Additionally, we propose an efficient iterative calibration process to optimize layer-specific decomposition by addressing the varying sensitivity of different LLM layers. ASVD can compress a network by 10-20%, without compromising the performance of LLMs. Based on the success of the low-rank decomposition of projection matrices in the self-attention module, we further introduce ASVD to compress the KV cache. By reducing the channel dimension of KV activations, memory requirements for KV cache can be largely reduced. Thanks to the 50-75% reduction in the rank of the KV projection matrices, ASVD can further achieve 50% KV cache reductions without performance drop in a training-free manner. Code and compressed models are available at ASVD4LLM.
@@ 1 Introduction
Numerous strategies have been proposed to address the memory consumption challenges in Large Language Models (LLMs), as reviewed by Zhu et al. [2023], Ding et al. [2023]. In the realm of LLM compression, various techniques have been extensively explored, including weight quantization [Dettmers et al., 2022], network pruning [Frantar & Alistarh, 2023], and knowledge distillation [Agarwal et al., 2023]. Distinct from these approaches, the paradigm of low-rank matrix decomposition is less explored but holds significant promise. Decomposition involves approximating the weight matrices in LLMs with matrices of lower rank, effectively reducing the model size. Given the massive number of parameters in LLMs, low-rank decomposition offers significant potential for memory reduction. Furthermore, low-rank decomposition can complement existing LLM compression techniques by further compressing quantized or pruned models, enhancing overall efficiency [Cheng et al., 2017]. Despite its potential, it remains a relatively underutilized approach for LLM compression.
@@
From the perspective of network compression, traditional low-rank decomposition methods typically adhere to a straightforward process: initially training the original model and subsequently fine-tuning the decomposed model [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021, Hsu et al., 2022]. While this approach is effective, it is resource-intensive and requires the entire training dataset and substantial computational power for end-to-end backpropagation. Applying this method to LLMs would encounter major challenges. Firstly, the training data for LLMs may not always be readily available, often restricted by privacy and commercial considerations. Secondly, the training process for these models is notoriously expensive, both in terms of time and computational resources. Given these constraints, the concept of “training-free” compression emerges as a more viable approach for LLMs [Zhu et al., 2023]. This approach includes methods like LLM post-training quantization [Dettmers et al., 2022, Yuan et al., 2023] and LLM post-training pruning [Frantar & Alistarh, 2023], which compress LLMs without the need for extensive retraining. These training-free (i.e., post-training) methods offer a more practical solution for efficiently compressing LLMs.
@@
To realize LLM low-rank decomposition in a training-free manner, we conduct an extensive analysis of the baseline methods for LLM decomposition. We first observe that straightforward application of existing low-rank decomposition techniques, which typically necessitate training, turns out ineffective for LLMs [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015, Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021]. Digging into the failures, we reveal several challenges unique to post-training decomposition for LLMs. One key challenge involves managing outliers in the activations, which significantly intensifies the decomposition error. This challenge echoes findings in LLM quantization research [Lin et al., 2023, Kim et al., 2023], highlighting the importance of handling such outliers in LLMs.
(a) Summarized Performance of ASVD.
(b) High-level idea of using ASVD to compress KV cache.
Figure 1: (a) Our post-training LLM decomposition method is orthogonal to existing LLM compression techniques, enabling it to function as a versatile and plug-and-play solution for prevalent compression paradigms, including popular quantization methods. (b) By applying low-rank decomposition via ASVD to the Key/Value projection matrices, the original high-dimensional KV cache can be replaced with a low-dimensional storage.
To address these challenges, we propose Activation-aware Singular Value Decomposition (ASVD), where the distribution of activations are considered into the weight decomposition process. Specifically, it achieves this by scaling the values in the weight matrix column-wisely, based on the distribution patterns observed across input and output channels. This adjustment proves particularly beneficial for channels with outlier activations, allowing the ASVD to allocate enhanced focus to these specific weights. Such targeted adjustment facilitates a more accurate reconstruction of the weight matrix, thereby minimizing prediction errors and improving the overall efficiency of the decomposition process. We further investigate the varying sensitivity of different LLM layers to decomposition. We find that weights in Multi-Head Attention layers  [Vaswani et al., 2017] tend to be more resilient to decomposition compared to those in Multi-Layer Perceptron layers. This sensitivity variability across layers prompts us to develop a method to assign the compression ratio for each layer. ASVD assesses each layer’s sensitivity to decomposition at different ranks, enabling us to assign a suitable rank for optimal decomposition. Note that this probing assess is very efficient, requiring only a limited sample set for evaluation.
Empirically, we evaluate ASVD with LLaMA and LLaMA-2 [Touvron et al., 2023a, b]. Our experiments reveal that ASVD can reduce the rank of the weight matrix by 50-75% (corresponds to 10-20% compression ratio), with only a marginal 1% accuracy loss on benchmarks like MMLU [Hendrycks et al., 2020]. We also validate ASVD is compatible with 4/6/8-bit quantization, as shown in Fig. 1(a).
Importantly, leveraging the successful low-rank decomposition of projection matrices in the self-attention module, we can integrate ASVD with KV cache compression. Specifically, by applying ASVD to decompose the Key/Value projection matrices, we can derive low-rank intermediate activations that serve as replacements for the KV cache stored in high-dimension space as shown in Fig. 1(b). This substitution significantly reduces the memory usage of the KV cache, enabling support for larger batch sizes or longer sequence lengths, which are essential for real-world applications [Yuan et al., 2024]. In practice, by replacing the KV cache with low-rank intermediate activations, we can reduce up to 50% of the memory consumption of KV cache.
@@ 2 Related Work
Large Language Model Compression. The field of model compression for Large Language Models (LLMs) has seen a surge of innovative techniques aimed at mitigating the substantial computation and memory requirements these models demand [Zhu et al., 2023, Yuan et al., 2024]. Various methods have emerged to address this challenge, each taking a unique approach to reduce the memory footprint of LLMs. These methods primarily fall into three categories: weight quantization [Courbariaux et al., 2015, Dettmers et al., 2022], network pruning [LeCun et al., 1989, Frantar & Alistarh, 2023], and knowledge distillation [Hinton et al., 2015, Agarwal et al., 2023]. For the wide body of research on LLM compression, please refer to [Zhu et al., 2023] for the comprehensive survey. Among these methods, weight quantization has gained significant traction in the context of LLMs due to its effectiveness. However, despite its popularity as a neural network compression technique, low-rank factorization has not been extensively explored in the realm of LLMs.
@@
Low-rank Decomposition. In the realm of low-rank decomposition [Schotthöfer et al., 2022] for neural network compression, existing methods can be broadly classified into two categories: fixed low rank and variable low rank approaches. Fixed rank methods typically involve decomposing weight matrices of pre-trained networks using techniques like Singular Value Decomposition (SVD) or tensor decomposition, followed by fine-tuning the factorized network [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015]. They also involve constraining weight matrices to maintain a fixed low rank during training [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021], or constructing layers as linear combinations of layers with varying ranks [Ioannou et al., 2015]. A notable limitation of these methods is the introduction of matrix decomposition rank as an additional hyperparameter requiring fine-tuning. In contrast, rank-adaptive methods address this limitation by automatically determining and adjusting the low-rank structure. In particular, Kim et al. [2015, 2019] apply heuristics search to pre-determine the decomposition rank, while Wen et al. [2017] learn low-rank weights through a loss function penalizing approximated matrix ranks. Li et al. [2023] use low-rank approximation plus a sparse matrix to compress the weight matrix in transformers.
@@
However, none of these methods have worked in the era of LLMs due to their training-require nature. In this work, we introduce ASVD, a post-training LLM decomposition approach enabling the adaptive determination of SVD ranks to optimize the matrix approximations based on feature activations. To our knowledge, this is the first attempt to compress the weights of LLMs through decomposition in a training-free manner.
@@ 3.1 Naïve SVD for Compressing Weight Matrix
Singular Value Decomposition (SVD) can be used to decomposing the weights of linear layers, which involves decomposing a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ into three matrices: $\mathbf{U}$, $\mathbf{\Sigma}$, and $\mathbf{V}^{T}$, such that $\mathbf{W}\approx\mathbf{W}\mathbf{\Sigma}\mathbf{V}^{T}$), where $\mathbf{\Sigma}$ is an $m\times n$ diagonal matrix, the diagonal values in $\mathbf{\Sigma}$ are the singular values of $\mathbf{W}$, and $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are corresponding right and left singular vector matrices, respectively [Demmel, 1997].
The SVD compression process for a weight matrix can be summarized in three steps: Decomposition: Factorize $\mathbf{W}$ using SVD. Truncation: Retain the top $k$ singular values and their corresponding right and left singular vectors. This results in approximated matrices $\mathbf{U}_{k}$, $\mathbf{\Sigma}_{k}$, and $\mathbf{V}_{k}^{T}$, where the right singular vector matrix $\mathbf{U}_{k}$ is $m\times k$, singular $\mathbf{\Sigma}_{k}$ is $k\times k$, and left singular vector matrix $\mathbf{V}_{k}^{T}$ is $k\times n$. The choice of $k$ is critical in balancing the compression ratio and the compressed model’s performance. Reconstruction: Reconstruct an approximated weight matrix: $\mathbf{W}_{k}=\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{U}_{k}^{T}$.
@@ 3.2 Challenges of Compressing LLMs via SVD
Decomposing the large matrices in LLMs (e.g., $4096\times 4096$ matrices ubiquitous in LLaMA-7b [Touvron et al., 2023a]) into lower ranks presents a viable pathway for model compression. However, straightforward application of existing low-rank decomposition techniques [Denton et al., 2014, Lebedev et al., 2014, Moczulski et al., 2015, Khodak et al., 2021, Wang et al., 2021, Li et al., 2023], which typically necessitate training, proves ineffective for LLMs.
Figure 2: Comparison between SVD and ASVD. Outlier channels in input activations ($\mathbf{X}$) are highlight in red, and ASVD takes these into consideration, which can contribute to a reduction in output error.
Challenge 1: Influence of Activation: This perspective shifts the focus from solely relying on the truncation error $\mathbf{L}_{\text{t}}$, which depends only on the model’s weights, to also accounting for the activations. The rationale behind this is the critical role of outliers in activations within LLMs  [Lin et al., 2023, Wei et al., 2022, Kim et al., 2023]. Thus, for effective LLM decomposition, our objective optimization becomes:
@@
To illustrate this concept, we visualize an example of $\mathbf{W}$, $\mathbf{W}_{k}$ (decomposed by simply SVD), $\mathbf{X}$, and the resulting variation in activations $\Delta\mathbf{Y}$ in Fig. 2 (Top line). This visualization reveals a critical insight: even when the variation in weights $\Delta\mathbf{W}=\mathbf{W}-\mathbf{W}_{k}$ is relatively minor, the corresponding variation in activations $\Delta\mathbf{Y}$ can be huge. This significant variation in activations is a key factor in why a straightforward SVD-based decomposition approach falls short in effectively decomposing LLMs. The activation variations, despite being derived from input activations of large magnitude (not the weight variations), can lead to considerable changes in the whole model’s output, thereby undermining the decomposition’s efficacy.
@@
These challenges underscore the necessity for innovative approaches specifically designed for the LLM decomposition. Our objective is to achieve efficient compression while circumventing the substantial computational and data demands associated with training-based methods. To address the first challenge, we introduce an Activation-aware SVD mechanism, which is detailed in Section 3.3. This method is designed to mitigate the impact of weight variation on activations. For the second challenge, we propose a Sensitivity-based Truncation Rank Searching mechanism, elaborated in Section 3.5, which adapts to the varying singular value distributions among different layers.
3.3 ASVD: Activation-aware Singular Value Decomposition
Our ASVD is designed to refine the weight matrix $\mathbf{W}$ in LLMs by taking into account the effect of input activation channels. The process comprises the following three steps: